OpenStudy (anonymous):

Find a polynomial function of degree 5. -1 as a zero multiplicity of 2. 3 as a zero multiplicity of 2. and 0 as a zero multiplicity of 1. Leading coefficient is 4.

OpenStudy (whpalmer4):

Okay, if you have a polynomial \(P(x)\) and you want to write it in factored form, with roots \(r_1, r_2, ...,r_n\), it can be written \[P(x) = a(x-r_1)(x-r_2)...(x-r_n)\]where \(a\) is a constant to arrange the polynomial to pass through a specified point, or have a given coefficient — it doesn't alter the roots, and can be chosen as 1 for convenience in many cases. If the multiplicity of a given root is > 1, then you need to raise that root's factor to its multiplicity as a power. For example, a root of 3 with multiplicity 2 will be \((x-3)^2\) in the factored polynomial.

hero (hero):

In other words, Compute f(x) = 4x(x + 1)^2(x - 3)^2

OpenStudy (whpalmer4):

Yeah, enjoy the algebra :-)

OpenStudy (anonymous):

So then 4(x^2+2x+1+x^2+6x+9)=0

OpenStudy (whpalmer4):

No, you have to multiply those together...

OpenStudy (whpalmer4):

\[4(x^2+2x+1)(x^2-6x+9) = \]

OpenStudy (anonymous):

How do you even foil that?..

OpenStudy (whpalmer4):

Not even that — we all dropped the \(x-0\) term! \[4x(x^2+2x+1)(x^2-6x+9)\]

OpenStudy (whpalmer4):

Distributive property...this is why I hate FOIL as a teaching method, no one knows what to do for bigger multiplications.

hero (hero):

I had it originally. I just copied someone else's mistake

hero (hero):

I don't know why they ever taught FOIL to begin with.

OpenStudy (whpalmer4):

:-) \[(a+b+c)(d+e+f) = a(d+e+f) + b(d+e+f) + c(d+e+f)\] don't forget to multiply that all by \(4x\) when you're done!

OpenStudy (whpalmer4):

Yeah, I don't understand that either, it's not like no one has ever seen the distributive property by the time they get to multiplying polynomials!

OpenStudy (whpalmer4):

Anyhow, @pita0001, as the problem says, you'll end up with a polynomial function of degree 5, which means you'll need to have an \(x^5\) term in there.

OpenStudy (whpalmer4):

These "construct a polynomial" problems do a bit of testing of understanding, but mostly they exercise your algebraic manipulation skills :-)

OpenStudy (anonymous):

Thank you both for your help..

OpenStudy (whpalmer4):

The only other bit I can think of that this problem didn't include is remembering that if you have a polynomial with only real coefficients, any complex roots must come in conjugate pairs, \(a\pm bi\). You'll undoubtedly see (if you haven't already) problems where they only give you one of the complex roots and you need to know how to construct the other.

OpenStudy (whpalmer4):

(and that you need to include it, of course)